OpenStudy (anonymous):

Solve the equation: a.) 4 cosx=1+2 cosx I have no idea what to do. Help!!!

OpenStudy (phi):

I would add -2 cos x to both sides (think of cos x as just another variable) 4 cos x - 2 cos x = 1 + 2 cos x - 2 cos x the 2 cos x - 2 cos x is 0 4 cos x - 2 cos x is 2 cos x. you get 2 cos x = 1 divide both sides by 2

OpenStudy (whpalmer4):

Sometimes to aid in thinking of cos x as just another variable, as phi suggests, it's easier to just rewrite the equation: 4c = 1 + 2c 2c = 1 c = 1/2 Now put the cos x back in cos x = 1/2